I am trying to figure out the balance of a college fund after 10 years if I invest $5000 at 3% compounded quarterly.
Found 2 solutions by stanbon, andkon:
Answer by stanbon(48526)

balance of a college fund after 10 years if I invest $5000 at 3% compounded quarterly.
-----------
A(t) = P(1/r/n)^(nt)
A(10) = 5000(1+0.03/4)^(4*10)
A(10) = 5000*(1.0075)^40
A(10) = 5000*1.348348612...
A(10) = $6741.74
==================
Cheers,
Stan H.

Answer by andkon(15)